# AnchorSeg: Language Grounded Query Banks for Reasoning Segmentation

### 20 Apr 2026 v1

@@ Abstract
Reasoning segmentation requires models to ground complex, implicit textual queries into precise pixel-level masks. Existing approaches rely on a single segmentation token <SEG>, whose hidden state implicitly encodes both semantic reasoning and spatial localization, limiting the model’s ability to explicitly disentangle what to segment from where to segment. We introduce AnchorSeg, which reformulates reasoning segmentation as a structured conditional generation process over image tokens, conditioned on language grounded query banks. Instead of compressing all semantic reasoning and spatial localization into a single embedding, AnchorSeg constructs an ordered sequence of query banks: latent reasoning tokens that capture intermediate semantic states, and a segmentation anchor token that provides explicit spatial grounding. We model spatial conditioning as a factorized distribution over image tokens, where the anchor query determines localization signals while contextual queries provide semantic modulation. To bridge token-level predictions and pixel-level supervision, we propose Token–Mask Cycle Consistency (TMCC), a bidirectional training objective that enforces alignment across resolutions. By explicitly decoupling spatial grounding from semantic reasoning through structured language grounded query banks, AnchorSeg achieves state-of-the-art results on ReasonSeg test set (67.7% gIoU and 68.1% cIoU). All code and models are publicly available at https://github.com/rui-qian/AnchorSeg.
@@ 4.1 Experimental Setting
Consistent with established methodologies in prior works (kazemzadeh2014referitgame; mao2016generation; lai2024lisa), we employ specific metrics for different tasks. For Reasoning Segmentation evaluation, we report gIoU and cIoU. Specifically, gIoU represents the mean of the Intersection-over-Union (IoU) scores calculated for each individual image, while cIoU is derived from the ratio of the cumulative intersection to the cumulative union across the dataset. For the Referring Expression Comprehension (REC) task, we adopt Precision@0.5 metric, where a prediction is considered correct if the IoU is at least 0.5.
@@ 5 Conclusion
Acknowledgments This work was supported by Dejing Dou’s Research Startup Fund from Fudan University and the computations in this research were performed using the CFFF platform of Fudan University.
6 Limitations
Although AnchorSeg effectively decouples spatial grounding from semantic reasoning through structured language grounded query banks, several limitations remain.
Dependence on LMM Quality. The quality of the language grounded query bank depends on the capability of the underlying LMM to generate meaningful latent reasoning tokens. Stronger LMMs produce more discriminative intermediate representations, which benefit our factorized formulation, while weaker models may yield less informative contextual queries and thus reduced performance gains. Future work could explore self-supervised objectives to improve the quality and structure of latent reasoning tokens.
Fixed Query Bank Size. AnchorSeg uses a fixed number $K$ of latent reasoning tokens for all queries, regardless of reasoning complexity. Simple queries (e.g., “the red car”) may not require multiple reasoning states, while highly complex queries (e.g., “the object that could prevent water damage in this scene”) might benefit from more intermediate reasoning states. Adaptive mechanisms that dynamically adjust the number of latent tokens based on query complexity could improve both efficiency and performance.
Computational Overhead. Compared to vanilla single-token approaches, AnchorSeg introduces additional computational costs through multiple query tokens, token-level similarity computation, and the TMCC training objective. While these components are essential for explicit disentanglement of semantic reasoning and spatial grounding, more efficient architectures or sparse attention mechanisms could reduce the computational burden for real-time applications.
Despite these limitations, AnchorSeg establishes a principled framework for structured conditional generation in reasoning segmentation, providing new insights into explicit language-vision alignment and factorized spatial reasoning.
@@ Appendix A Ethics and Societal Impact
Table 9: Overview of the training dataset mixture.
@@
Pascal Part chen2014detect 4.37K n.a. n.a. n.a.
@@
reason_seg_plus instance_seg yang2023lisa++ 58.50K n.a. n.a. n.a.
@@ Appendix B Datasets
Following the data organization paradigm established by prior works, we construct a comprehensive training mixture to equip our model with multi-granularity visual understanding and complex reasoning capabilities. As shown in Table  9, our training data is categorized into eight primary types: (1) sem_seg for fundamental semantic understanding; (2) refer_seg for standard referring expression comprehension; (3) neg_refer_seg and (4) correct_refer_seg for robust handling of false premises; (5) vqa for general conversational abilities; and finally, (6) reason_seg, (7) reason_seg_plus, and (8) multi_reason_seg to facilitate advanced logical deduction and multi-target grounding.
@@
Semantic Segmentation (sem_seg). This category incorporates widely-used public datasets, including ADE20K, COCO-Stuff, Pascal Part, PACO-LVIS, and Mapillary. To align with the instruction-following objective, we randomly select specific categories for each image and reformulate the data into a visual question-answering (VQA) format using predefined templates.
@@
Referring Segmentation (refer_seg). The RefClef dataset (also referred to as ReferItGame) is a foundational benchmark for referring expression comprehension, comprising 130K expressions for nearly 100K regions across 20,000 images. Unlike later COCO-based benchmarks, RefClef includes annotations for both discrete object instances and amorphous “stuff” categories.
@@
The RefCOCO series (including RefCOCO, RefCOCO+, and RefCOCOg) consists of large-scale benchmarks for referring expression comprehension and segmentation based on MS-COCO. RefCOCO and RefCOCO+ contain concise phrases, with the latter prohibiting absolute spatial terms to emphasize appearance attributes. In contrast, RefCOCOg features longer, more complex descriptive sentences.
@@
RefZom is a challenging benchmark established to address the “one-to-one” assumption limitation in traditional referring segmentation by incorporating three distinct settings: one-to-zero (no target), one-to-one (single target), and one-to-many (multiple targets). It comprises a total of 55,078 images and 74,942 annotated objects. The dataset is partitioned into a training set of 43,749 images (58,356 objects) and a testing set of 11,329 images (16,586 objects).
@@
G-RefCOCO is a large-scale benchmark for Generalized Referring Expression Segmentation (GRES), comprising 278,232 expressions, including 80,022 multi-target and 32,202 empty-target ones, referring to the objects in 19,994 images. The images are split into four subsets: training, validation, test-A, and test-B, following the same UNC partition of RefCOCO.
@@
Negative Referring Segmentation (neg_refer_seg). R-RefCOCO(+/g) datasets extend RefCOCO, RefCOCO+, and RefCOCOg by introducing negative text inputs that require empty mask predictions. Adhering to the original UNC and UMD partitions, they augment the training data with a 1:1 positive-negative ratio and significantly expand the validation sets with approximately 10 negative descriptions per reference object to rigorously evaluate robustness.
@@
Correct Referring Segmentation (correct_refer_seg). FP-RefCOCO(+/g) extend the RefCOCO series to facilitate the False Premise Correction task. The datasets utilize large language models to create context-aware false premises by modifying specific objects, attributes, or relations. Preserving original splits, the benchmark maintains a 1:1 positive-to-negative ratio across roughly 20K-26K images per dataset to evaluate reasoning robustness.
@@
Visual Question Answering (vqa). We utilize the LLaVA-Instruct-150K dataset, which consists of approximately 157.71K vision-language instruction samples. This dataset is incorporated to maintain the model’s general conversational proficiency and its ability to follow diverse natural language instructions.
@@
Reasoning Segmentation (reason_seg). ReasonSeg is a benchmark established to evaluate reasoning segmentation capabilities, comprising 1,218 image-instruction-mask data samples. The images are annotated with implicit text instructions that necessitate complex reasoning or world knowledge, categorized into short phrases and long sentences. To ensure reliable assessment, the dataset is partitioned into 239 training, 200 validation, and 779 testing samples, allocating a larger proportion to testing for rigorous evaluation.
Enhanced Reasoning Segmentation (reason_seg_plus). ReasonSeg-Plus is based on the COCO2017 dataset. The Instance Segmentation subset focuses on training models for instance-level segmentation, while the CoT (Chain-of-Thought) subset enhances global scene understanding and reasoning segmentation. Additionally, the Conversation subset equips models with the ability to perform segmentation within dynamic dialogues, and the Caption subset integrates segmentation tasks directly into image captioning workflows.
@@
Multi-target Reasoning Segmentation (multi_reason_seg). MUSE is a large-scale benchmark established for the Multi-target Reasoning Segmentation task, comprising 246K question-answer pairs and 910K instance-level mask annotations sourced from the LVIS dataset. It features complex queries with arbitrary numbers of open-set targets, averaging 3.7 objects per response. The dataset is partitioned into 239K training, 2.8K validation, and 4.3K testing samples to rigorously evaluate pixel-level reasoning and understanding capabilities.
@@ Appendix C Implementation Details
Given an input image $\mathbf{x}_{img}\in\mathbb{R}^{h\times w\times c}$ and a textual prompt $\mathbf{x}_{txt}$, LLaVA ($\mathbf{x}_{img}$, $\mathbf{x}_{txt}$) produces the image token embedding $\boldsymbol{I}_{img}\in\mathbb{R}^{576\times 4096}$ and the <SEG> token embedding $\boldsymbol{q}_{anc}\in\mathbb{R}^{4096}$. The similarity map is computed as $\boldsymbol{q}_{anc}\times\boldsymbol{I}_{img}^{T}=(1\times 4096)\times(576\times 4096)^{T}=1\times 576$, followed by reshaping ($576\rightarrow 24\times 24$) into $1\times 24\times 24$, and then interpolating ($1\times 24\times 24\rightarrow 1\times 256\times 256$) and convolving ($1\times 256\times 256\rightarrow 256\times 64\times 64$) to obtain the spatial prior $\mathbf{P}\in\mathbb{R}^{256\times 64\times 64}$. Note that SAM’s dense visual features are given by $\mathbf{f}=\mathcal{G}_{\mathcal{V}}^{enc}(\mathbf{x}_{img})\in\mathbb{R}^{256\times 64\times 64}$. We inject the spatial prior $\mathbf{P}$ into the visual feature map $\mathbf{f}$ via element-wise addition, yielding conditioned visual features $\tilde{\mathbf{f}}=\mathbf{f}\oplus\mathbf{P}$ for SAM mask generation.
@@
C.1 Interpolation Details for Token–Mask Cycle Consistency (TMCC)
We specify the interpolation operators used in TMCC and refer to Algorithm 1 for the complete token-to-image resizing pipeline.
Token-to-Mask interpolation.
@@ C.1 Interpolation Details for Token–Mask Cycle Consistency (TMCC)
For the reverse direction $\mathcal{L}_{M2T}$, the ground-truth mask $\mathbf{M}\in\{0,1\}^{H\times W}$ is first mapped to the $L_{vl}=336$ canvas using the same resize-and-pad operator as above (bilinear resize with align_corners=False, antialias=True, followed by bottom/right zero padding), and then downsampled to the token grid using nearest-neighbor interpolation:
$\mathbf{M}^{grid}$ $=\mathrm{Interp}\!\left(\mathbf{M}^{336};\,G\times G,\,\texttt{nearest}\right),$ (13)
$\mathbf{M}^{grid}$ $\in\{0,1\}^{G\times G}.$
@@
$\mathbf{M}^{grid}_{\sigma}$ $=\mathbf{M}^{grid}*G_{\sigma}\in[0,1]^{G\times G},$ (14)
$\mathbf{M}^{\downarrow}_{\sigma}$ $=\mathrm{vec}\!\left(\mathbf{M}^{grid}_{\sigma}\right)\in[0,1]^{N}.$
@@ C.2 Extending the Anchor Query to Multiple Spatial Anchors
Multi-anchor variants introduce additional computational overhead, which scales approximately linearly with $T$ when decoding is repeated, and can lead to anchor competition and duplicate predictions. To address these issues, practical strategies include anchor-wise NMS/merging, diversity regularization across anchors, and bipartite matching-based assignment during training when instance-level supervision is available. We leave the design of more efficient and principled multi-anchor generation and training strategies for future work.

### 22 Apr 2026 v3

@@ Abstract
Reasoning segmentation requires models to ground complex, implicit textual queries into precise pixel-level masks. Existing approaches rely on a single segmentation token <SEG>, whose hidden state implicitly encodes both semantic reasoning and spatial localization, limiting the model’s ability to explicitly disentangle what to segment from where to segment. We introduce AnchorSeg, which reformulates reasoning segmentation as a structured conditional generation process over image tokens, conditioned on language grounded query banks. Instead of compressing both semantic reasoning and spatial localization into a single embedding, AnchorSeg constructs an ordered sequence of query banks: latent reasoning tokens that capture intermediate semantic states, and a segmentation anchor token that provides explicit spatial grounding. We model spatial conditioning as a factorized distribution over image tokens, where the anchor query determines localization signals while contextual queries provide semantic modulation. To bridge token-level predictions and pixel-level supervision, we propose Token–Mask Cycle Consistency (TMCC), a bidirectional training objective that enforces alignment across resolutions. By explicitly decoupling spatial grounding from semantic reasoning through structured language grounded query banks, AnchorSeg achieves state-of-the-art results on ReasonSeg test set (67.7% gIoU and 68.1% cIoU). All code and models are publicly available at https://github.com/rui-qian/AnchorSeg.
@@ 4.1 Experimental Setting
Consistent with established evaluation protocols in prior works (kazemzadeh2014referitgame; mao2016generation; lai2024lisa), we adopt task-specific metrics. For reasoning segmentation, we report gIoU and cIoU. Specifically, gIoU is the mean Intersection-over-Union (IoU) computed per image, while cIoU is defined as the ratio of the cumulative intersection to the cumulative union over the dataset. For referring expression comprehension (REC), we use Precision@0.5, where a prediction is considered correct if its IoU with the ground truth is at least 0.5.
@@ 5 Conclusion
Acknowledgments
This work was supported by Dejing Dou’s Research Startup Fund from Fudan University and the computations in this research were performed using the CFFF platform of Fudan University.
Limitations
Fixed Query Bank Size. AnchorSeg uses a fixed number $K$ of latent reasoning tokens for all queries, regardless of reasoning complexity. However, simple queries (e.g., “the red car”) do not necessarily require multiple reasoning states, whereas more complex queries (e.g., “the object that could prevent water damage in this scene”) often benefit from richer intermediate reasoning. This mismatch suggests that adaptive mechanisms that dynamically allocate the number of latent tokens based on query complexity could improve both computational efficiency and overall performance.
@@ Appendix A Ethics and Societal Impact
Table 9: Overview of the training data mixture and its composition across different task categories.
@@
Pascal-Part chen2014detect 4.37K n.a. n.a. n.a.
@@
reason_seg_plus InstanceSeg yang2023lisa++ 58.50K n.a. n.a. n.a.
@@ Appendix B Datasets
Following prior data organization paradigms, we construct a comprehensive training mixture to support multi-granularity visual understanding and complex reasoning. In Table 9, we organize the training data into eight primary types: (1) sem_seg for fundamental semantic understanding; (2) refer_seg for standard referring expression comprehension; (3) neg_refer_seg and (4) correct_refer_seg for robust handling of false premises; (5) vqa for general conversational ability; and (6) reason_seg, (7) reason_seg_plus, and (8) multi_reason_seg for advanced reasoning and multi-target grounding.
@@
Semantic Segmentation (sem_seg). This category includes several widely used datasets, such as ADE20K, COCO-Stuff, Pascal-Part, PACO-LVIS, and Mapillary. To align with the instruction-following objective, we randomly select target categories for each image and reformulate them as visual question-answering (VQA) pairs using predefined templates.
@@
Referring Segmentation (refer_seg). The RefCLEF dataset (a.k.a. ReferItGame) is a widely used benchmark for referring expression comprehension, containing 130K expressions over nearly 100K regions across 20K images. In contrast to later COCO-based benchmarks, RefCLEF includes annotations for both discrete object instances and amorphous “stuff” categories.
@@
The RefCOCO series (including RefCOCO, RefCOCO+, and RefCOCOg) comprises several large-scale benchmarks for referring expression comprehension and segmentation based on MS COCO. RefCOCO and RefCOCO+ contain relatively concise phrases, with RefCOCO+ further restricting absolute spatial terms to emphasize appearance attributes. In contrast, RefCOCOg features longer and more complex descriptions.
@@
RefZOM is a challenging benchmark designed to address the one-to-one assumption in traditional referring segmentation, covering three settings: one-to-zero (no target), one-to-one (single target), and one-to-many (multiple targets). It contains 55,078 images and 74,942 annotated objects, split into 43,749 training images (58,356 objects) and 11,329 testing images (16,586 objects).
@@
gRefCOCO is a large-scale benchmark for Generalized Referring Expression Segmentation (GRES), comprising 278,232 expressions, including 80,022 multi-target and 32,202 empty-target cases, over 19,994 images. The dataset is split into four subsets—training, validation, test-A, and test-B—following the UNC split of RefCOCO.
@@
Negative Referring Segmentation (neg_refer_seg). The R-RefCOCO, R-RefCOCO+, and R-RefCOCOg datasets extend RefCOCO, RefCOCO+, and RefCOCOg by introducing negative text inputs that require empty-mask predictions. Following the original UNC and UMD splits, they augment the training data with a 1:1 positive-to-negative ratio and expand the validation sets with approximately 10 negative descriptions per reference object to evaluate robustness.
@@
Correct Referring Segmentation (correct_refer_seg). FP-RefCOCO(+/g) extend the RefCOCO series for the false premise correction task. They use large language models to generate context-aware false premises by modifying objects, attributes, or relations. Following the original splits, each dataset maintains a 1:1 positive-to-negative ratio with approximately 20K–26K images to evaluate reasoning robustness.
@@
Visual Question Answering (vqa). We use the LLaVA-Instruct-150K dataset, which contains approximately 158K vision-language instruction samples. It is included to maintain the model’s general conversational ability and its capability to follow diverse natural language instructions.
@@
Reasoning Segmentation (reason_seg). ReasonSeg is a benchmark for reasoning segmentation, containing 1,218 image-instruction-mask samples. The images are annotated with implicit text instructions requiring complex reasoning or world knowledge, categorized into short phrases and long sentences. The dataset is split into 239 training, 200 validation, and 779 test samples.
Reasoning Segmentation Plus (reason_seg_plus). ReasonSeg-Plus is built on COCO 2017 and consists of four subsets: Instance Segmentation for instance-level segmentation, CoT (Chain-of-Thought) for global scene understanding and reasoning segmentation, Conversation for segmentation in multi-turn dialogue, and Caption for integrating segmentation with image captioning.
@@
Multi-target Reasoning Segmentation (multi_reason_seg). MUSE is a large-scale benchmark for multi-target reasoning segmentation, comprising 246K question-answer pairs and 910K instance-level mask annotations built on LVIS. The dataset features complex queries with arbitrary numbers of open-set targets, averaging 3.7 objects per response. It is split into 239K training, 2.8K validation, and 4.3K test samples to evaluate pixel-level reasoning and understanding.
@@ Appendix C Implementation Details
Given an input image $\mathbf{x}_{img}\in\mathbb{R}^{h\times w\times c}$ and a textual prompt $\mathbf{x}_{txt}$, LLaVA produces the image token embeddings $\boldsymbol{I}_{img}\in\mathbb{R}^{576\times 4096}$ and the <SEG> token embedding $\boldsymbol{q}_{anc}\in\mathbb{R}^{4096}$. The similarity map is computed as
$\boldsymbol{q}_{anc}\boldsymbol{I}_{img}^{T}\in\mathbb{R}^{1\times 576},$
which is then reshaped to $1\times 24\times 24$, interpolated to $1\times 256\times 256$, and convolved to produce a spatial prior $\mathbf{P}\in\mathbb{R}^{256\times 64\times 64}$. The dense visual features from SAM are given by $\mathbf{f}=\mathcal{G}_{\mathcal{V}}^{enc}(\mathbf{x}_{img})\in\mathbb{R}^{256\times 64\times 64}$. We inject the spatial prior via element-wise addition, yielding conditioned features $\tilde{\mathbf{f}}=\mathbf{f}\oplus\mathbf{P}$ for mask generation. The training recipes of AnchorSeg across different benchmarks are summarized in Table 10.
C.1 Interpolation Details for Token–Mask Cycle Consistency (TMCC)
We specify the interpolation operators used in TMCC and refer to Algorithm (1) for the complete token-to-image resizing pipeline.
Token-to-Mask Interpolation.
@@ C.1 Interpolation Details for Token–Mask Cycle Consistency (TMCC)
For the reverse direction $\mathcal{L}_{M2T}$, the ground-truth mask $\mathbf{M}\in\{0,1\}^{H\times W}$ is first mapped to the $L_{vl}=336$ canvas using the same resize-and-pad operator as above (bilinear resize with align_corners=False, antialias=True, followed by bottom/right zero padding), and then downsampled to the token grid using nearest-neighbor interpolation as
$\begin{aligned} \mathbf{M}^{grid}&=\mathrm{Interp}\!\left(\mathbf{M}^{336};\,G\times G,\,\texttt{nearest}\right),\\ \mathbf{M}^{grid}&\in\{0,1\}^{G\times G}.\end{aligned}$
@@
$\begin{aligned} \mathbf{M}^{grid}_{\sigma}&=\mathbf{M}^{grid}*G_{\sigma}\in[0,1]^{G\times G},\\ \mathbf{M}^{\downarrow}_{\sigma}&=\mathrm{vec}\!\left(\mathbf{M}^{grid}_{\sigma}\right)\in[0,1]^{N}.\end{aligned}$
@@ C.2 Extending the Anchor Query to Multiple Spatial Anchors
Multi-anchor variants introduce additional computational overhead, which scales approximately linearly with $T$ when decoding is repeated. They may also lead to anchor competition and duplicate predictions. To mitigate these issues, practical strategies include anchor-wise non-maximum suppression (NMS) or merging to remove duplicate outputs, diversity regularization across anchors, and bipartite matching-based assignment during training when instance-level supervision is available. We leave the design of more efficient and principled multi-anchor generation and training strategies to future work.